# A. Abed Abud CERN, The European Organization for Nuclear Research, 1211 Meyrin, Switzerland B. Abi

### v1

@@ III.1 Neutrino-Electron Elastic Scattering
Figure 2: Comparison of the cross-section-weighted energy and angular distributions of electrons for the CC events in red and the eES events in blue. The angular plots are shown as a function of the cosine of the angle between the neutrino direction and the final-state electron direction. For both types of interactions, the supernova energy spectrum is assumed for the incoming neutrinos. Top left: distribution of the energy of incoming neutrinos (continuous line) and outgoing electrons (dashed line). Top right: distribution of the cosine of the scattering angle. Bottom: two-dimensional distributions showing the relative number of events as a function of both the neutrino energy and the cosine of the scattering angle, for the two types of interactions.
@@
The primary observable of this interaction is the emitted ; additional observables are the de-excitation products of the excited potassium nucleus in the final state. The output energy and angular distributions of this interaction are also shown in Fig. 2 in red and Fig. 1 (a) depicts an example from Geant4.
@@
Events from CC interactions are considerably more abundant than the aforementioned eES events (about 3000 events at a core collapse distance of 10 kpc for DUNE compared to 300 for eES in Tab. 1). Yet, the direction of the electron trajectory correlates very weakly with the neutrino direction, due to the two competing nuclear transition types, Fermi and Gamow-Teller. The electrons produced by CC-induced interactions governed by Fermi transitions have an angular distribution described by , while Gamow-Teller transitions correspond to [31] with respect to the neutrino direction. For our assumed supernova neutrino energy spectrum, the angle dependence of the two contributions happen to cancel each other out almost exactly, resulting in a nearly isotropic angular distribution for CC events, as shown in the red distribution in Fig. 2(b). Figure 3 depicts the contribution of the two matrix elements for the GKVM supernova spectrum.
The pointing resolution for a supernova burst therefore depends on a precise classifier between eES and CC events as discussed in Sec. V.
We note that the Fermi/Gamow-Teller cancellation is not perfect for other assumed flux spectra (e.g. Ref. [31]), resulting in a weak anisotropy of the CC-absorption-induced electrons for many cases. Furthermore, forbidden transitions, not currently included in MARLEY, have a backward angular distribution [32] which may have an effect on the overall CC anisotropy. Currently, there are large (and not fully understood) uncertainties on the relative components of CC nuclear transition type and hence on the CC angular distribution (as well as on the total cross section [33].)
@@
We note that while the expected CC anisotropy is relatively weak, it should be possible to extract directional information from the CC events, especially if the competing nuclear transition types can be tagged and examined separately. In principle, the nuclear transition types can be distinguished statistically using their different nuclear de-excitation product distributions. However, given the large uncertainties on the CC angular distribution555Note that these uncertainties may be addressed directly with independent laboratory measurements. and that our assumed spectrum produces the most experimentally challenging situation— a directional signal on top of a uniform CC distribution— we evaluate the pointing ability under this simple assumption and leave detailed study and use of CC anisotropies for the future.
Figure 3: The contribution of Fermi and Gamow-Teller transitions to the angular distribution for CC events, generated from MARLEY for the GVKM flux model. For this model’s neutrino spectrum, the angular correlation cancels out to a good approximation.
@@ IV.1 Simulation of events in DUNE
The eES and CC event generators in LArSoft produce particles uniformly in the workspace geometry with random neutrino directions sampled from an isotropic angular distribution. Standard radioactive and detector noise models were used during the simulations [15]. The considered radioactive background sources include contaminants of the liquid argon such as 39Ar and 85Kr, radioactive contaminants from the TPC, and neutrons from the cavern walls. We expect radiological and cosmogenic background to have a relatively small effect on the pointing quality. The dominant 39Ar background of 1 Bq/kg results in only of the order of one background-induced blip expected in the spatial vicinity of a supernova event over one drift time [34]. Therefore we expect our evaluation to be robust against evolutions of the DUNE background model. Furthermore, the background will be known precisely; it can be fully characterized from the data taken near in time to the supernova burst, which will enable optimization of background mitigation for reconstruction algorithms. The simulated electronics signal process accounts for the charge deposition and drift physics, the field response of the sensing wires, electronics and digitization response of the front-end electronics, as well as the inherent electronics noise according to Ref. [34]. The Projection Matching Algorithm [35] is used for 3D track reconstruction.
An example eES event display is shown in Fig. 4. Of note, the longest track marked by ”0” (several tens of cm) with high charge deposition corresponds to the primary electron and there are as well as several shorter tracks. These tracks result from lower energy final-state particles that were produced from the interaction (such as the deexcitation products of CC events), or secondary particles created as the primary electron travels in the detector. Notably, some secondary particles have directions that correlate to the primary electron, making them useful in determining the starting direction of the primary electron track (discussed in Sec. IV.2). Additionally, charge may also be registered as a result of background radioactive decay events, such as the decay of and electronics noise. These signals are very short compared to the primary electron, are far apart from each other, and are of low amplitude, which are all characteristics that can be used for extracting signals from background.
Figure 4: Example standard DUNE event display of an eES event. The colors show a heat map of the charge deposited on each wire segment per time tick. From top to bottom, the three plots show the view of the three wire planes, U, V, and X respectively, for one APA of the LArTPC. The trajectories with numerical labels are reconstructed tracks, where “0” corresponds to the primary electron track, while ”1” and ”2” refer to the brems particles.
@@ IV.2 Primary track identification and energy reconstruction
The reconstruction algorithm consists of several steps.
First, the energy depositions associated with the supernova neutrinos are identified. The first step in reconstructing the event is to determine which reconstructed track corresponds to the primary electron (marked with “0” in Fig. 4). The DUNE version of the projection matching algorithm as described in [36] is applied for the track reconstruction. We show in the following that the overall performance of the low-energy reconstruction is sufficient for the task at hand. Specific mitigations for the intrinsic head-tail ambiguity of the track will be described. In a noise- and background-free scenario, the track of the primary electron typically has the greatest length (see Fig. 1), as the electron deposits more charge compared to its secondary particles. However, the presence of radioactive decay particles can complicate this, as a highly energetic radioactive decay product may deposit more charge than the primary electron and create a longer track. These scenarios are rejected by examining the spatial distance between tracks. Particles associated with the neutrino interactions are clustered together, while the radioactive particles are spaced out over the total simulated LAr volume. Therefore, the primary track is selected by examining the relative position of charge depositions. Reconstructed 3D space points are sorted by their corresponding hit charges, and for the ten space points with the highest charge depositions, the distances between these space points are calculated. The two space points with the closest spatial distance between each other among those ten space points are picked and one of them is selected as the estimated reconstructed vertex position. While this location may not exactly be the true vertex, it allows the rough localization of the position inside the detector. The track with the most associated charge in the proximity of this reconstructed vertex (within a sphere of radius of 120 cm) is finally selected to be the primary electron track.
@@ IV.3 Direction reconstruction and head-tail disambiguation
Most relevant for the supernova pointing is the primary electron’s direction. The goal is to attain a accurate “reconstruction resolution”, which goes along with minimizing the sky area that must be surveyed to confidently encompass the supernova. We expect a head-tail ambiguity in LArTPC track directions as the charge drift speed towards the anode is slow compared to the propagation of particles through the TPC material. Consequently, the track’s head-tail direction cannot be determined through timing. In other words, it is not known through timing which side of the track corresponds to the origin of the track, causing the reconstructed direction to potentially be opposite of the true charged-particle track direction. This results in a bimodal distribution of angle between the true and reconstructed directon around the true and reversed directions of the primary track. The reconstruction algorithm aims to enhance the peak pointing towards the true direction.
@@
To evaluate the performance of the reconstruction resolution, we consider detection probabilities from scanning the sky regions with the highest likelihood values first. The cosine angles between the reconstructed direction and the truth direction are collected into a histogram. The histogram is integrated from both the forward and backward directions inwards, up until the histogram drops below a certain value. This value is lowered until the integral equals 68% of the total integral of the distribution. The resolution metric “sky fraction” is defined to be the fraction of the sky that is covered by the integration. The sky fraction can be used to define an equivalent angular resolution by multiplying with 4. If the two sides of the distribution are close to fully symmetric, the sky fraction would be twice as large as the fully unambigiuous case, corresponding to an equal probability of success for finding the supernova on each side of the sky.
@@
The head-tail ambiguity can be resolved statistically by looking at the adjacent tracks created by secondary particles. The bremsstrahlung gamma rays from the primary electron subsequently release electrons in argon atoms via Compton scattering. These secondary electrons correlate with the forward direction of the primary electron. The directional correlation is closer when the primary electron is of higher energy. The particular procedure of disambiguating the direction of the primary electron is carried out in a process called “brems flipping”.
To apply the method, a starting-point and an end-point of the primary electron track are assumed arbitrarily. The vectors from the starting point as well as from the end point of the primary track to each secondary track are determined. Subsequently, the cosine value of the angle between each vector pointing to the secondary tracks and the vector along the primary track is evaluated. The average of these cosine values is calculated.
In the next step, the assignment of the starting-point and the end-point are switched and the same calculations are carried out. Of the two vertices, the one with the larger average cosine value is finally selected as the actual starting point of the primary track. Most of the secondary particles are emitted towards the end of the primary track in a forward direction, leading to a preference for larger cosine values (corresponding to smaller angles) as can be understood from Fig. 5. The brems-flipping method is also described in Algorithm 1 (see Appendix).
Figure 5: Illustration of brems flipping: The angles between the brems particles (blue) and the primary track marked in black correspond to the actual direction of the primary electron, while the ones marked in red belong to the incorrect opposite direction. The average of the cosine of these angles is larger in the case of the correct set of angles.
@@ IV.4 Performance of the reconstruction algorithm on single events
The performance of the reconstruction algorithm before the application of brems flipping is shown in the top plot of Fig. 6 in blue. The distribution is centered around both the parallel and anti-parallel directions with respect to the true direction, with a minor preference for the parallel direction. The reason for this preference is that the projection matching algorithm creates many track candidates at first and then merges candidates that are close to each other and have a small angle between them. If a secondary track is sufficiently close to the primary track, the merging process may combine it with the primary track, flipping the primary track to the correct orientation.
Brems flipping extends the same principle to the separately detected brems particles, meaning that their location with respect to the track indicates the direction of the primary track.
By applying brems flipping, the magnitude of the distribution in red in the anti-parallel direction is significantly decreased, confirming that the technique is a valuable tool in resolving the ambiguity in track directions.
@@
The bottom plot of Fig. 6 shows how brems flipping reduces the sky fraction for mono-energetic electron energies. The electron energy spectrum from eES events is given as a reference for the energy regime of interest for supernova pointing. Brems flipping has a higher impact at higher primary electron energies, as those result in more secondary tracks. The successful application of brems flipping on the full supernova neutrino MC is presented in Sec. V.
@@
Our studies show that the performance of the tracking algorithm as a function of the initial track direction with respect to the detector coordinate system is not perfectly isotropic, as expected. The particle track’s inclination with respect to the readout wire planes affects the charge deposition on the wires [35] and therefore influences the uncertainty of the reconstructed direction. The worst performance is observed along the drift direction (detector coordinate ).
Figure 7(a) demonstrates this anisotropy. For this figure, it is assumed that there are no directional ambiguities, to reveal the full impact of this anisotropy.
Figure 7(b) is created assuming the actual performance of the reconstruction algorithm including brems flipping for the head-tail disambiguation.
While the reconstruction performance varies with track direction, the figures show that the track direction ambiguity has a stronger detector anisotropy. We therefore first evaluate pointing capability assuming a uniform performance in directional reconstruction. Subsequently, we investigate the performance of the method as a function of supernova direction in detector coordinates.
Figure 6: Effectiveness of the brems-flipping algorithm: The top plot shows the bimodal distribution of the angular difference between true and reconstructed electron directions using the supernova eES electron energy spectrum, centering around the parallel () and anti-parallel () directions. Brems flipping significantly decreases the magnitude of the anti-parallel peak. The bottom plot shows the relationship between the covered sky fraction and mono-energetic electron energy. The black curve corresponds to a perfect directional disambiguation always resulting the true direction. Brems flipping performs better at higher energies, as there are more secondary tracks to reference. The electron energy spectrum from eES in gray illustrates the energies relevant to supernova pointing.
Figure 7: Pointing resolution, defined in terms of “sky fraction” as described in the text, of individual electron events as a function of the electron’s true direction in the detector coordinate system to study anisotropies in the performance. The coordinate system is defined with being the drift directions, and being approximately the beam direction. (shown vertically) and (shown horizontally) are spherical coordinates in a coordinate system for which correspond to the direction and corresponds to the direction. The supernova eES electron energy distribution is used.
(a) shows the pointing resolution given perfect track head-tail disambiguation (b) depicts the pointing resolution for the actual performance of the reconstruction algorithm including the brems-flipping algorithm.
@@ V Maximum likelihood method for burst pointing
A maximum likelihood method [13] is used to reconstruct the direction of a supernova burst ensemble of events from the reconstructed information of individual neutrino events. For an event with a known interaction type, the probability density function (PDF) has the functional form of , where the index indicates the neutrino channel: eES or CC. and are the reconstructed energy and direction of the specific event, and is the direction of the supernova. The PDF is formulated to be a function of the reconstructed energy as well as the inner product of the supernova direction and reconstructed electron direction, , as we make the aforementioned assumption that the event direction reconstruction quality is uniform as a function of the true neutrino direction666In principle residual detector response anisotropy can be taken into account in the likelihood PDFs.. The PDFs are normalized by energy bin, as the electron energy distribution does not affect the direction likelihood. LArSoft simulations of one million events in each channel are used to determine the PDFs. The Monte Carlo samples are divided into energy bins, spanning from 0 to 40 MeV with a width of 2 MeV per bin, from 40 MeV to 70 MeV at 10 MeV per bin, and a final bin spanning from 70 MeV from 100 MeV. Varying energy bin widths are used to ensure sufficient counting statistics at high energies, where the expected numbers of events are low.
@@
Figure 8: PDFs of eES and CC events. The eES PDF demonstrates a high correlation between the supernova and primary electrons’ direction increasing towards higher electron energies. The CC PDF is mostly flat, as the correlation between primary electron and neutrino directions is weak.Figure 9: Several bins for the PDF of the eES interaction in Figure 8, re-binned to 10 MeV / bin. The relative magnitude of the peak near (the parallel direction) increases as energy increases because brems flipping improves at high energies. The width of the true direction peaks also decreases, indicating a decrease in directional variance.
@@
Here, a perfect classification between eES and CC events is assumed, . The effect of imperfect classifications is discussed later. By parameterizing using the azimuth and zenith angles, the supernova direction is reconstructed via a two-variable minimization. For the minimization, events that have reconstructed electron energies of less than 5 MeV are excluded, as they are found to reduce pointing performance due to the relatively large uncertainties in their energy and direction reconstruction. The minimization is done by a grid search. The grid search provides the likelihood function value for all directions in the sky, which is illustrated in a directional map.
The reconstructed neutrino directions associated with a typical supernova burst are shown in Fig. 10, and the output of the reconstruction procedure is shown in Fig. 11. Two local minima are 180 degrees away from each other due to the remaining ambiguity in track direction.
Figure 10: An example directional map filled with the reconstructed electron directions for a simulated supernova burst, eES events carrying the directional information are marked in blue. Event statistics are shown for a core collapse at a distance of 10 kpc, and 40 kton of fiducial mass in the detector.Figure 11: An example directional map filled with the negative log-likelihood values and confidence contours. This map is computed from the same burst shown as in Fig. 10. Reconstruction is done assuming the classification parameters of and and a successful direction reconstruction is achieved with the actual supernova direction marked with a star in the figure.

### v2

@@ III.2 Electron-Neutrino Charged-Current Absorption Interactions
The primary observable of this interaction is the emitted ; additional observables are the de-excitation products of the excited potassium nucleus in the final state. The output energy and angular distributions of this interaction are also shown in Fig. 2 (a) and (b) in red and Fig. 1 (a) depicts an example CC event from Geant4.
@@
Events from CC interactions are considerably more abundant than the aforementioned eES events (about 3000 events at a core collapse distance of 10 kpc for DUNE compared to 300 for eES in Tab. 1.) Yet, the direction of the electron trajectory correlates very weakly with the neutrino direction, due to the two competing nuclear transition types, Fermi and Gamow-Teller. The electrons produced by CC-induced interactions governed by Fermi transitions have an angular distribution described by , while Gamow-Teller transitions correspond to [gardiner2021nuclear] with respect to the neutrino direction. For our assumed supernova neutrino energy spectrum and cross-section model, the angular dependences of the two contributions happen to cancel each other out almost exactly, resulting in a nearly isotropic angular distribution for CC events, as shown in the red distribution in Fig. 2(b). Figure 3 depicts the contribution of the two matrix elements for the GKVM supernova spectrum. The pointing resolution for a supernova burst therefore depends on a precise classifier between eES and CC events as discussed in Sec. V.
@@
We note that the Fermi/Gamow-Teller cancellation is not perfect for other assumed flux spectra (e.g., Ref. [gardiner2021nuclear]), resulting in a weak anisotropy of the CC-absorption-induced electrons for many cases. Furthermore, forbidden transitions, not currently included in MARLEY, have a backward angular distribution [VanDessel:2019obk] which may have an effect on the overall CC anisotropy. Currently, there are large (and not fully understood) uncertainties on the relative components of CC nuclear transition type and hence on the CC angular distribution (as well as on the total cross section [DUNE:2023rtr].) We note that while the expected CC anisotropy is relatively weak, it should be possible to extract directional information from the CC events, especially if the competing nuclear transition types can be tagged and examined separately. In principle, the nuclear transition types can be distinguished statistically using their different nuclear de-excitation product patterns555This approach is under study for future application.. However, given the large uncertainties on the CC angular distribution666Note that these uncertainties may be addressed directly with independent laboratory measurements. and that our assumed spectrum produces the most experimentally challenging situation— a directional signal on top of a uniform CC distribution— we evaluate the pointing ability under this simple assumption and leave detailed study and use of CC anisotropies for the future.
@@ IV.1 Simulation of events in DUNE
The eES and CC event generators in LArSoft produce particles uniformly in the workspace geometry with random neutrino directions sampled from an isotropic angular distribution. Standard radioactive and detector noise models were used during the simulations [abi2020deep]. The considered radioactive background sources include contaminants of the liquid argon such as 39Ar and 85Kr, radioactive contaminants from the TPC, and neutrons from the cavern walls. We expect radiological and cosmogenic background to have a relatively small effect on the pointing quality. The dominant 39Ar background of 1 Bq/kg results in only of the order of one background-induced blip expected in the spatial vicinity of a supernova event over one drift time [abi2020volume]. Therefore, we expect our evaluation to be robust against evolutions of the DUNE background model. Furthermore, the background will be known precisely; it can be fully characterized from the data taken near in time to the supernova burst, which will enable optimization of background mitigation for reconstruction algorithms. The simulated electronics signal process accounts for the charge deposition and drift physics, the field response of the sensing wires, electronics and digitization response of the front-end electronics, as well as the inherent electronics noise according to Ref. [abi2020volume]. The Projection Matching Algorithm [Antonello:2012hu] is used for 3D track reconstruction.
An example eES event display is shown in Fig. 4. Of note, the longest track marked by “0” (several tens of cm) with high charge deposition corresponds to the primary electron and there are as well as several shorter tracks. These tracks result from lower energy final-state particles that were produced from the interaction (such as the deexcitation products of CC events), or secondary particles created as the primary electron travels in the detector. Notably, some secondary particles have directions that correlate to the primary electron, making them useful in determining the starting direction of the primary electron track (discussed in Sec. IV.2.) Additionally, charge may also be registered as a result of background radioactive decay events, such as the decay of and electronics noise. These signals are very short compared to the primary electron, are far apart from each other, and are of low amplitude, which are all characteristics that can be used for extracting signals from background.
@@ IV.2 Primary track identification and energy reconstruction
The reconstruction algorithm consists of several steps. First, the energy depositions associated with the supernova neutrinos are identified. The first step in reconstructing the event is to determine which reconstructed track corresponds to the primary electron (marked with “0” in Fig. 4.) The DUNE version of the projection matching algorithm as described in [Antonello:2012hu] is applied for the track reconstruction. We show in the following that the overall performance of the low-energy reconstruction is sufficient for the task at hand. Specific mitigations for the intrinsic head-tail ambiguity of the track will be described. In a noise- and background-free scenario, the track of the primary electron typically has the greatest length (see Fig. 1,) as the electron deposits more charge compared to its secondary particles. However, the presence of radioactive decay particles can complicate this, as a highly energetic radioactive decay product may deposit more charge than the primary electron and create a longer track. These scenarios are rejected by examining the spatial distance between tracks. Particles associated with the neutrino interactions are clustered together, while the radioactive particles are spaced out over the total simulated LAr volume. Therefore, the primary track is selected by examining the relative position of charge depositions. Reconstructed 3D space points are sorted by their corresponding hit charges, and for the ten space points with the highest charge depositions, the distances between these space points are calculated. The two space points with the closest spatial distance between each other among those ten space points are picked and one of them is selected as the estimated reconstructed vertex position. While this location may not exactly be the true vertex, it allows the rough localization of the position inside the detector. The track with the most associated charge in the proximity of this reconstructed vertex (within a sphere of radius of 120 cm) is finally selected to be the primary electron track.
@@ IV.3 Direction reconstruction and head-tail disambiguation
Most relevant for the supernova pointing is the primary electron’s direction. The goal is to attain an accurate “reconstruction resolution”, which goes along with minimizing the sky area that must be surveyed to confidently encompass the supernova. We expect a head-tail ambiguity in LArTPC track directions as the charge drift speed towards the anode is slow compared to the propagation of particles through the TPC material. Consequently, the track’s head-tail direction cannot be determined through timing. In other words, it is not known through timing which side of the track corresponds to the origin of the track, causing the reconstructed direction to potentially be opposite of the true charged-particle track direction. This results in a bimodal distribution of angle between the true and reconstructed directon around the true and reversed directions of the primary track. The reconstruction algorithm aims to enhance the peak pointing towards the true direction.
@@
To evaluate the performance of the reconstruction resolution, we consider detection probabilities from scanning the sky regions with the highest likelihood values first. A histogram is filled with the distribution of cosine angles between the reconstructed direction and the truth direction. The histogram is integrated from both the forward and backward directions inwards for all probability densities exceeding a threshold; the threshold is lowered until the integral equals 68% of the total integral of the distribution. The resolution metric “sky fraction” is defined to be the fraction of the sky that is covered by the 68% integration. Calculation of the sky fraction is graphically illustrated in Fig. 5. The sky fraction converts to an equivalent area of coverage in solid angle by multiplying by . For a small-angle approximation, the solid angle corresponds to times the opening angle squared, which can be considered the pointing resolution. Similarly, multiplying the sky fraction by two yields an effective confidence interval in . If the two sides of the distribution are close to fully symmetric, the sky fraction would be twice as large as the fully unambiguous case, corresponding to an equal probability of success for finding the supernova on each side of the sky.
@@
The head-tail ambiguity can be resolved statistically by looking at the adjacent tracks created by secondary particles. The bremsstrahlung gamma rays from the primary electron subsequently release electrons in argon atoms via Compton scattering. These secondary electrons correlate with the forward direction of the primary electron. The directional correlation is closer when the primary electron is of higher energy. The particular procedure of disambiguating the direction of the primary electron is carried out in a process called “brems flipping”. To apply the method, a starting-point and an end-point of the primary electron track are assumed arbitrarily. The vectors from the starting point as well as from the end point of the primary track to each secondary track are determined. Subsequently, the cosine value of the angle between each vector pointing to the secondary tracks and the vector along the primary track is evaluated. The average of these cosine values is calculated. In the next step, the assignment of the starting-point and the end-point are switched and the same calculations are carried out. Of the two vertices, the one with the larger average cosine value is finally selected as the actual starting point of the primary track. Most of the secondary particles are emitted towards the end of the primary track in a forward direction, leading to a preference for larger cosine values (corresponding to smaller angles) as can be understood from Fig. 6. The brems-flipping method is also described in Algorithm 1 (see Appendix).
@@ IV.4 Performance of the reconstruction algorithm on single events
The performance of the reconstruction algorithm before the application of brems flipping is shown in the top plot of Fig. 7 in blue. The distribution is centered around both the parallel and anti-parallel directions with respect to the true direction, with a minor preference for the parallel direction. The reason for this preference is that the projection matching algorithm creates many track candidates at first and then merges candidates that are close to each other and have a small angle between them. If a secondary track is sufficiently close to the primary track, the merging process may combine it with the primary track, flipping the primary track to the correct orientation. Brems flipping extends the same principle to the separately detected brems particles, meaning that their location with respect to the track indicates the direction of the primary track. By applying brems flipping, the magnitude of the distribution in red in the anti-parallel direction is noticeably decreased, confirming that the technique is a valuable tool in resolving the ambiguity in track directions.
@@
The bottom plot of Fig. 7 shows how brems flipping reduces the sky fraction for mono-energetic electron energies. The electron energy spectrum from eES events is given as a reference for the energy regime of interest for supernova pointing. Brems flipping has a higher impact at higher primary electron energies, as those result in more secondary tracks. This fact motivates us to include energy reconstruction as part of the supernova direction reconstruction, as much can be gained by weighing higher-energy events more than lower-energy ones. The successful application of brems flipping on the full supernova neutrino MC is presented in Sec. V.
@@
Our studies show that the performance of the tracking algorithm as a function of the initial track direction with respect to the detector coordinate system is not perfectly isotropic, as expected. The particle track’s inclination with respect to the readout wire planes affects the charge deposition on the wires [Antonello:2012hu] and therefore influences the uncertainty of the reconstructed direction. The worst performance is observed along the drift direction (detector coordinate ) due to projection of the track over fewer wires. Figure 8(a) demonstrates this anisotropy. For this figure, it is assumed that there are no directional ambiguities, to reveal the full impact of this anisotropy. Figure 8(b) is created assuming the actual performance of the reconstruction algorithm including brems flipping for the head-tail disambiguation. While the reconstruction performance varies with track direction, the figures show that the track direction ambiguity has a stronger detector-related anisotropy. We therefore first evaluate pointing capability assuming a uniform performance in directional reconstruction. Subsequently, we investigate the performance of the method as a function of supernova direction in detector coordinates.
@@ V Maximum likelihood method for burst pointing
A maximum likelihood method [abe2016real] is used to reconstruct the direction of a supernova burst ensemble of events from the reconstructed information of individual neutrino events. For an event with a known interaction type, the probability density function (PDF) has the functional form of , where the index indicates the neutrino channel: eES or CC. and are the reconstructed energy and direction of the specific event, and is the direction of the supernova. The PDF is formulated to be a function of the reconstructed energy as well as the inner product of the supernova direction and reconstructed electron direction, , as we make the aforementioned assumption that the event direction reconstruction quality is uniform as a function of the true neutrino direction777In principle, residual detector response anisotropy can be taken into account in the likelihood PDFs.. The PDFs are normalized by energy bin, as the electron energy distribution does not affect the direction likelihood. LArSoft simulations of one million events in each channel are used to determine the PDFs. The Monte Carlo samples are divided into energy bins, spanning from 0 to 40 MeV with a width of 2 MeV per bin, from 40 MeV to 70 MeV at 10 MeV per bin, and a final bin spanning from 70 MeV from 100 MeV. Varying energy bin widths are used to ensure sufficient counting statistics at high energies, where the expected numbers of events are low.
@@
Here, a perfect classification between eES and CC events is assumed, . The effect of imperfect classifications is discussed later. By parameterizing using the azimuth and zenith angles, the supernova direction is reconstructed via a two-variable minimization. For the minimization, events that have reconstructed electron energies of less than 5 MeV are excluded, as they are found to degrade pointing performance due to the relatively large uncertainties in their energy and direction reconstruction. The minimization is done by an adaptive grid search: after the initial coarse search of the entire phase space we applay a second grid search of a smaller area. The result of the second grid search is further refined using a local minimization of the PDF initialized at the fitted grid point. The reconstructed neutrino directions associated with a typical supernova burst are shown in Fig. 11, and the negative-log likelihoods values for all possible supernova directions for this burst are shown in Fig. 12. Two local minima are 180 degrees away from each other due to the remaining ambiguity in track direction.